y=-3/5x+1
if x=-5
how do i work the problem to find y

Answer by FightinBlueHens(27)

Assuming 5x is in the denominator: -3/(5x):
1. Substitute -5 for x in the problem: y= -3/(5*-5) + 1
2. Simplify the denominator: y= -3/-25 + 1
3. Simplify the fraction by getting rid of extra symbols: since the negative sign is both on the top and the bottom, it cancels out, because a negative divided by a negative is a positive so the fraction is 3/25. If you convert that to a decimal that's .12
4. Add 1 to the fraction or decimal you now have: .12 + 1 = 1.12
If you are working with a fraction, convert the 1 to a fraction that equals 1 and that has the same denominator as the other fraction (25).
Remember, that any number divided by itself equals 1 so change the 1 to 25/25 and add it to 3/25, which gives you 28/25.
Another way to do it is to say that 1 + 3/25 or 1 And 3/25 is a complex fraction, and you want to convert it to an imporper fraction to get one number. To do that multiply the integer (in this case 1) by the denominator, 25, and then add 3 to it. This gives you the numerator, 28, and we leave the denominator as 25.
If 1 = 25/25, do you see why we multiply 1 by 25 and then add 3 in the second way?